0 0
Add a comment Improve this question Transcribed image text
Answer #1

First you have to get water's density in order to find it's volume and the length of the water column in the right arm.

d = density of water
m= mass
v = volume

d = m/v

so

v = m/d = 100gr/(1g/cm^3) = 100cm^3

now you know that

a= area
h = length

v = a*h

so

h = v/a = 100cm^3/4.6cm^2 = 21.74 cm

that would be item a

now for b you know that the pressures in the contact area of water and mercury have to be the same, so

P1 = pressure of water

P2 = pressure of mercury

d1 = density of water

d2 = density of mercury

g = acceleration of gravity

h1 = height of water

h2 = height of mercury

P1 = P2

d1gh1=d2gh2

d1h1=d2h2

so

h2 = d1h1/d2

h2 = (1gm/cm^3)(21.74cm)/(13.6gm/cm^3)= 1.5985cm

so that would be the height of mercury from the point it touches the water.

Hope this helps.
Answer #2

First you have to get water's density in order to find it's volume and the length of the water column in the right arm.

d = density of water
m= mass
v = volume

d = m/v

so

v = m/d = 100gr/(1g/cm^3) = 100cm^3

now you know that

a= area
h = length

v = a*h

so

1) h = v/a = 100cm^3/4.6cm^2 = 21.74 cm

that would be item a

now for b you know that the pressures in the contact area of water and mercury have to be the same, so

P1 = pressure of water

P2 = pressure of mercury

d1 = density of water

d2 = density of mercury

g = acceleration of gravity

h1 = height of water

h2 = height of mercury

P1 = P2

d1gh1=d2gh2

d1h1=d2h2

so

2) h2 = d1h1/d2

h2 = (1gm/cm^3)(21.74cm)/(13.6gm/cm^3)= 1.5985cm

so that would be the height of mercury from the point it touches the water.

Hope this helps.
Answer #3

First you have to get water's density in order to find it's volume and the length of the water column in the right arm.

d = density of water
m= mass
v = volume

d = m/v

so

v = m/d = 100gr/(1g/cm^3) = 100cm^3

now you know that

a= area
h = length

v = a*h

so

1) h = v/a = (100cm^3)/(4.6cm^2) = 21.74 cm

that would be item a

now for b you know that the pressures in the contact area of water and mercury have to be the same, so

P1 = pressure of water

P2 = pressure of mercury

d1 = density of water

d2 = density of mercury

g = acceleration of gravity

h1 = height of water

h2 = height of mercury

P1 = P2

d1gh1=d2gh2

d1h1=d2h2

so

2) h2 = d1h1/d2

h2 = (1gm/cm^3)(21.74cm)/(13.6gm/cm^3)= 1.5985cm

so that would be the height of mercury from the point it touches the water.

Hope this helps
Answer #4

1) 21.74 cm
2) 0.4968 cm